# Simultaneous false discovery bounds for invariant causal prediction

### v2

@@ A.3 Proof of Proposition 3
When $H_{0,S^{*}}(\mathcal{E})$ is not rejected, for any $i\not\in S^{*}$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}>\alpha$, so $i\not\in\widehat{S}(\mathcal{E})$. Hence, we have $\widehat{S}(\mathcal{E})\subseteq S^{*}$. Therefore,
@@ A.4 Proof of Proposition 4
When $H^{*}_{0,S}$ is true, we have $S^{*}\subseteq[m]\setminus S$ as $S\cap S^{*}=\emptyset$, so $p^{*}_{S}=\max_{I\subseteq[m]\setminus S}p_{I}\geq p_{S^{*}}$. Therefore, for any $c\in(0,1)$,
@@
A.5 Proof of Theorem 1
Theorem A.1.
Let $\alpha\in(0,1)$, then
$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$
@@ A.5 Proof of Theorem 1
For any $R\subseteq[m]$,
$p^{*}_{R\cap\mathcal{N}^{*}}=\max_{I\subseteq[m]\setminus(R\cap\mathcal{N}^{*})}p_{I}\geq\max_{I\subseteq[m]\setminus\mathcal{N}^{*}}p_{I}=p^{*}_{\mathcal{N}^{*}}.$
So if $p^{*}_{\mathcal{N}^{*}}>\alpha$, we have $p^{*}_{R\cap\mathcal{N}^{*}}>\alpha$, which implies that $|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)$ by definition. Hence
$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq P(p^{*}_{\mathcal{N}^{*}}>\alpha).$
@@
A.6 Proof of Proposition 5
Proposition A.5.
Let $S^{*}$ be the set of true causal predictors. Consider the desired case where only $H_{0,S^{*}}(\mathcal{E})$ is true, $p_{S^{*}}\sim U(0,1)$ and $p_{S}=0$ for any $S\neq S^{*}$. Then, for $p^{\prime}_{i}$ satisfying (15), there exists some $c^{*}\in(0,1)$ such that $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})>c^{*}$.
@@ A.6 Proof of Proposition 5
For any $i$ such that $H^{*}_{0,i}$ is true, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}=p_{S^{*}}$. Since $P_{H^{*}_{0,i}}(p^{\prime}_{i}<p^{*}_{i})>0$, there must exist some rational number $c^{*}\in(0,1)$ such that
$P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*}<p^{*}_{i})>0.$ (16)
Then,
$P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})$
$=$ $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
$=$ $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i}\leq c^{*})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
$>$ $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*},p^{\prime}_{i}<p^{*}_{i})+P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*},p^{\prime}_{i}=p^{*}_{i})$
$=$ $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c^{*})$
$=$ $P_{H^{*}_{0,i}}(p_{S^{*}}\leq c^{*})$
$=$ $c^{*},$
which completes the proof. ∎

### v3

@@ Appendix B Proofs
(Proposition 3.3) When $H_{0,S^{*}}(\mathcal{E})$ is not rejected, for any $i\not\in S^{*}$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}>\alpha$, so $i\not\in\widehat{S}(\mathcal{E})$. Hence, we have $\widehat{S}(\mathcal{E})\subseteq S^{*}$. Therefore,
@@
(Proposition 4.1) When $H^{*}_{0,S}$ is true, we have $S^{*}\subseteq[m]\setminus S$ as $S\cap S^{*}=\emptyset$, so $p^{*}_{S}=\max_{I\subseteq[m]\setminus S}p_{I}\geq p_{S^{*}}$. Therefore, for any $c\in(0,1)$,
@@
B.5 Proof of Theorem 4.1
This result directly follows from the fact that $t_{\alpha}(R)$ is equivalent to the simultaneous true discovery bound based on closed testing, as discussed in Section 4.2. We also give a self-contained proof below for completeness.
@@
(Theorem 4.1) If $p^{*}_{\mathcal{N}^{*}}>\alpha$, we have $|R\cap S^{*}|=|R\setminus\mathcal{N}^{*}|\geq t_{\alpha}(R)$ by definition.
Hence
$P(|R\cap S^{*}|\geq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq P(p^{*}_{\mathcal{N}^{*}}>\alpha).$
@@
B.6 Proof of Theorem 4.2
@@
(Theorem 4.2) Let
$u_{\alpha}(R)=|R|-t_{\alpha}(R)=\max\{|I|:I\subseteq R\text{ and}p^{*}_{I}>\alpha\}.$
and
$\tilde{u}_{\alpha}(R)=|R|-\tilde{t}_{\alpha}(R)=\underset{S\subseteq[m]\text{ with }p_{S}>\alpha}{\max}|R\setminus S|$
Then, to show that $t_{\alpha}(R)=\tilde{t}_{\alpha}(R)$, it suffices to show that $u_{\alpha}(R)=\tilde{u}_{\alpha}(R)$.
Let
$A(R)=\{I\subseteq R:\max_{U\subseteq[m]\setminus I}p_{U}>\alpha\}\quad\text{and}\quad B=\{S\subseteq[m]:p_{S}>\alpha\},$
then we have
$u_{\alpha}(R)=\max\{|I|:I\subseteq R\text{ and }p^{*}_{I}>\alpha\}=\max\{|I|:I\in A(R)\}$
and
$\tilde{u}_{\alpha}(R)=\max\{|R\setminus S|:S\subseteq[m]\text{ and }p_{S}>\alpha\}=\max\{|R\setminus S|:S\in B\}.$
Let $I_{1}\in A(R)$ and $S_{2}\in B$ be the sets for which $u_{\alpha}(R)$ and $\tilde{u}_{\alpha}(R)$ take their maximal values, respectively. That is,
$u_{\alpha}(R)=|I_{1}|\quad\text{and}\quad\tilde{u}_{\alpha}(R)=|R\setminus S_{2}|=:|I_{2}|.$
First, note that $I_{2}=R\setminus S_{2}\subseteq R$ implies that $S_{2}\subseteq[m]\setminus I_{2}$, so $\underset{U\subseteq[m]\setminus I_{2}}{\max}p_{U}\geq P_{S_{2}}>\alpha$, where the second inequality holds as $S_{2}\in B$. Hence, we have $I_{2}\in A(R)$, which implies that $|I_{2}|\leq|I_{1}|$ by the definition of $I_{1}$.
To show that $u_{\alpha}(R)=\tilde{u}_{\alpha}(R)$, it is sufficient to prove that $|I_{1}|\leq|I_{2}|$.
Let $S_{1}=\operatorname*{argmax}_{U\subseteq[m]\setminus I_{1}}p_{U}$, so $S_{1}\subseteq[m]\setminus I_{1}$. Together with the fact that $I_{1}\subseteq R$, we have $I_{1}\subseteq R\setminus S_{1}$, so $|I_{1}|\leq|R\setminus S_{1}|$. In addition, since $I_{1}\in A(R)$, we have $p_{S_{1}}>\alpha$. That is, $S_{1}\in B$. Hence, $|R\setminus S_{1}|\leq|R\setminus S_{2}|$ by the definition of $S_{2}$. Finally, we have $|I_{1}|\leq|R\setminus S_{1}|\leq|R\setminus S_{2}|=|I_{2}|$. ∎